# Navigating the Path of Writing: Outline-guided Text Generation with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have significantly impacted the writing process, enabling collaborative content creation and enhancing productivity. However, generating high-quality, user-aligned text remains challenging. In this paper, we propose Writing Path, a framework that uses explicit outlines to guide LLMs in generating goal-oriented, high-quality pieces of writing. Our approach draws inspiration from structured writing planning and reasoning paths, focusing on capturing and reflecting user intentions throughout the writing process. We construct a diverse dataset from unstructured blog posts to benchmark writing performance and introduce a comprehensive evaluation framework assessing the quality of outlines and generated texts. Our evaluations with GPT-3.5-turbo, GPT-4, and HyperCLOVA X demonstrate that the Writing Path approach significantly enhances text quality according to both LLMs and human evaluations. This study highlights the potential of integrating writing-specific techniques into LLMs to enhance their ability to meet the diverse writing needs of users.
@@ 1 Introduction
Writing is a fundamental means of structuring thoughts and conveying knowledge and personal opinions Collins and Gentner (1980). This process requires systematic planning and reviewing in detail. Hayes (1980) describes writing as a complex problem-solving process and explores how planning and execution interact in writing. That is, writing involves more than merely generating text; it encompasses developing a proper understanding of the topic, gathering relevant subject matter, and implementing thorough structuring.
@@
Figure 1: Comparative overview of writing approaches: (A) direct generation with an LM, (B) iterative LM-based writing involving planning, editing, and explaining, and (C) the proposed Writing Path method, which starts with a consistency-focused plan, incorporates information-rich browsing, and results in an augmented, consistent, and rich outline.
@@
Recent advancements in Large Language Models (LLMs) have significantly impacted the writing workflow, enhancing both its efficiency and productivity. One significant area of exploration is the collaborative use of LLMs in writing processes Lee et al. (2022); Mysore et al. (2023), which typically involves the establishment of a writing plan and iterative improvement of interim outputs through revision Schick et al. (2023); Yang et al. (2022), as illustrated in Figure 1 (b). Here, the focus is on leveraging the remarkable generative capabilities of LLMs to gain advantages in terms of fluency, consistency, and grammatical accuracy. Productivity tools like Notion AI, Jasper, Writesonic, sudoWrite, and Cohesive are practical examples of leveraging LLMs to support users in creating content more efficiently. However, generating high-quality content that accurately aligns with specific user intentions and requirements still holds room for improvement Wang et al. (2024).
@@
To address this, we propose Writing Path, a methodology designed to incorporate user requirements such as desired topic, textual flow, keyword inclusion, and search result integration into the writing process. Writing Path emphasizes the importance of systematic planning and a clear outline from the early stages of writing. Inspired by the structured writing plan of Hayes (1980) and the reasoning path of Wei et al. (2022), the Writing Path collects ideas and creates outlines that encapsulate the user’s intentions prior to producing the specific content rather than directly outputting the final text. Furthermore, the initial outlines are further augmented with additional information through information browsing. Such a structured approach offers enhanced control over the text generation process and improves the quality of the content produced by LLMs.
@@
We also utilize a multi-aspect writing evaluation framework to assess the intermediate and final productions from the Writing Path, offering a way to evaluate the quality of free-form text  [^1] Free-form text generation focuses on creating diverse texts tailored to specific information and user intentions, unlike story generation, which develops narratives with cohesive plots and characters without relying on reference texts. Taking into account that conventional Likert scales (1-5 point ratings) Clark ; Hinkin (1998) are inappropriate for evaluating creative tasks, particularly those related to writing Chakrabarty et al. (2023), our evaluation framework aims to provide more precise and reliable assessments for the outlines and final texts.
For evaluation purposes, we construct a free-form blog text dataset incorporating a wide range of writing styles and topics from real users, including Beauty, Travel, Gardening, Cooking, and IT. Using this dataset, we evaluate how well the LLM outputs reflect the user’s intentions. Applying the Writing Path to various LLMs, including GPT-3.5-turbo Brown et al. (2020), GPT-4 Achiam et al. (2023), and HyperCLOVA X Yoo et al. (2024), shows significant performance gains across all evaluated models. These results validate that our approach enables the models to maintain a stronger focus on the given topic and purpose, ultimately generating higher-quality text that more accurately reflects user intentions.
@@
2 Related Work
2.1 Collaborative Writing with Language Models
Recursive Reprompting and Revision technique extends the potential of LMs to assist not only with draft generation but also with editing and revision processes. This approach employs LLM prompt chains such as planning - drafting - reviewing - suggesting revisions in an iterative fashion to enhance the quality of written content. Re3 (Yang et al., 2022) introduces a framework for maintaining the long-range coherence of draft generation. It operates separate rewriter and edit modules in its prompt chain to check and refine plot relevance and long-term factual consistency. PEER (Schick et al., 2023) proposes a recursive revision framework based on the concept of self-training, where the model autonomously selects the editing operations for revision and provides explanations for the modifications it makes. RECURRENTGPT Zhou et al. (2023) utilizes a recursive, language-based mechanism to simulate LSTM Hochreiter and Schmidhuber (1997), enabling the generation of coherent and extended texts.
2.2 Integrating External Information
2.3 Writing Evaluation
Figure 2: Main architecture of Writing Path, our proposed framework for guiding large language models to generate high-quality text following a structured writing process. The Writing Path condenses text generation into five key steps. This approach emulates the human writing planning process, enabling the model to produce text that aligns with the specified writing goals.
@@ 3 Design of Writing Path
We propose Writing Path, a systematic writing process to produce consistent, rich, and well-organized text with LLMs. Our method is inspired by how humans write; as humans typically collect materials and write an outline before writing the full text, Writing Path consists of five steps (Figure 2): (1) prepare metadata, (2) generate title and initial outline, (3) browse for information, (4) generate augmented outline, and (5) write the text. Each step is guided by a specific prompt configuration that aligns LLM output with specific step requirements. We also utilize information browsing to gather relevant content and keywords tailored to the writing stage.
The key steps of Writing Path are those that generate outlines as they establish a structured writing plan. Research suggests that a well-structured outline significantly impacts the quality of the written text Collins and Gentner (1980); Sun et al. (2022); Yang et al. (2022, 2023). The initial sketch is transformed into a detailed outline, including the flow, style, keywords, and relevant information from search results. This outline provides a clearer view of the final text to the LLMs, enabling them to generate a draft that better aligns with the intent of the user. The specific steps are described as follows:
Step 1: Prepare Meta Data
The first step establishes the writing direction and target reader using metadata $m$, which includes i) purpose, ii) type, iii) style, and iv) keywords. To simulate this process, we converted human-written texts into metadata (see Section 5.1 for details of the dataset).
Step 2: Generate Title and Initial Outline
The second step generates the title $t$ and initial outline $O_{\text{init}}$ based on the metadata $m$ from step 1, using the LLM function $f_{\text{llm}}$ with a prompt configuration function $\phi_{s}$. Here, $s$ indicates the step index, and for step 2, the prompt configuration is $\phi_{2}$:
$t,O_{\text{init}}=f_{\text{llm}}(\phi_{2}(m)),$ (1)
@@
Step 3: Browse for Information
The third step enriches the text by collecting additional information and keywords to reinforce the initial outline. We use the search function $f_{\text{search}}$ with the generated title $t$ as the query to retrieve the top-1 blog document, $D_{\text{sim}}$:
$D_{\text{sim}}=f_{\text{search}}(t)$ (2)
In our implementation, we employ the NAVER search API [^2] https://developers.naver.com/docs/serviceapi/ search/blog/blog.md to retrieve the top-1 document among similar blog posts. From the blog document, we extract keywords $K$ using the $f_{\text{llm}}$ with a prompt configuration $\phi_{3}$:
$K=f_{\text{llm}}(\phi_{3}(D_{\text{sim}})),$ (3)
The extracted keywords and additional information from the search results lead to a more specific writing plan, improving the quality of the generated text.
Step 4: Generate Augmented Outline
The fourth step refines the initial outline by adding subheadings and specific details to each section based on incorporating the keywords collected from the previous step. The augmented outline $O_{\text{aug}}$ is generated using the LLM function $f_{\text{llm}}$ with a prompt configuration $\phi_{4}$ that takes the title $t$, keywords $K$, and initial outline $O_{\text{init}}$ as inputs:
@@
The resulting augmented outline, $O_{\text{aug}}$, comprises headers ($h_{i,0}$) and their corresponding subheaders ($h_{i,j}$), where $i$ denotes the header index, and $j$ indexes the subheaders. This outline with detailed structure serves as a comprehensive writing plan, breaking down the text into manageable parts and providing clear direction for the content.
Step 5: Write the Text
Finally, the text for each section $d^{i}$ is generated using the LLM function with a prompt configuration $\phi_{5}$ that takes the title $t$ and the corresponding section of the augmented outline $O_{\text{aug}}^{i}$ as inputs:
$d^{i}=f_{\text{llm}}(\phi_{5}(t,O_{\text{aug}}^{i}))$ (5)
@@
$D=\{d^{1},d^{2},\ldots,d^{n}\}$ (6)
Writing Path organically connects each step of the writing process, employing an outline to aggregate and manage complex and diverse information, and assists users in producing high-quality writing. The prompts utilized for the Writing Path are detailed in the Appendix 9.
@@ 4 Evaluation of Writing Path
Evaluating the effectiveness of Writing Path compared to existing writing support systems is challenging. Most previous studies do not directly utilize outlines in the writing process and, therefore, lack systematic methods to assess outline quality. Even when outlines are used, evaluation relies only on human evaluation Yang et al. (2023); Zhou et al. (2023). Moreover, the heavy reliance on human evaluation in current approaches is also problematic when assessing full texts Schick et al. (2023); Yang et al. (2022); Lee et al. (2023); Fabbri et al. (2021), as it requires evaluating various aspects of the written work.
To address these limitations, we propose a novel evaluation framework for Writing Path that combines human and automatic evaluation to assess the quality of generated outlines and final texts from multiple perspectives. The proposed method establishes criteria for evaluating outline quality and sets standards for assessing the quality of final texts, enabling objective and reproducible validation of Writing Path’s effectiveness as a writing support system.
@@ 4.1.1 Automatic Evaluation
We adapt various metrics to evaluate the logical alignment, coherence, diversity, and repetition in outlines, following criteria established in linguistic literature Van Dijk (1977); Pitler and Nenkova (2008); Tang et al. (2019); Elazar et al. (2021).
  • Logical alignment: Based on Chen and Eger (2023), we utilize Natural Language Inference (NLI) which examines whether the headers and subheaders within an outline logically connect, ensuring the structural integrity necessary for coherent argumentation333we utilize gpt-4-turbo for NLI evaluation.
Note that coherency and diversity exhibit a trade-off relationship; maintaining coherence while covering a wide range of topics is essential to ensure the effectiveness of the outline in guiding the writing process.
@@ 4.1.2 Human Evaluation
In addition to automatic evaluation metrics, we conduct a human evaluation to assess aspects of the generated outlines that are difficult to capture solely with automatic measures. The human evaluation criteria are based on aspects considered in previous studies on text coherence, relevance, and quality assessment Yang et al. (2022, 2023); Zhou et al. (2023); Ke et al. (2022). For both initial and augmented outlines, the human evaluation is performed on the following five aspects, using a 1-4 point scale:
  • Natural Flow: Assesses whether the outline flows in a natural order.
4.2 Writing Evaluation
Traditional evaluation metrics such as Likert scales (1-5 point ratings) are not well-suited for assessing creative tasks like long story generation Chakrabarty et al. (2023). Acknowledging the need for more specific writing evaluation methods, we employ CheckEval Lee et al. (2024) to assess writing quality.
The Lee et al. (2024) reports an average Kendall Tau correlation of 0.4925 between human and LLM evaluations for the summarization task, exhibiting a higher level of agreement compared to other evaluation methods. This suggests that CheckEval effectively captures writing quality, establishing it as an appropriate tool for assessing the quality of model-generated texts.
CheckEval decomposes the evaluation aspects into more granular sub-questions, forming a detailed checklist. These aspect-based checklists can make performance evaluations by either humans or LLMs more fine-grained. Moreover, by explicitly capturing the evaluator’s reasoning behind each rating, this approach enhances the explainability of the evaluation process. The CheckEval paper reports an average Kendall Tau correlation of 0.4925 between human and LLM evaluations for the summarization task, exhibiting a higher level of agreement compared to other evaluation methods. This suggests that CheckEval effectively captures writing quality, establishing it as an appropriate tool for assessing the quality of model-generated texts. To adapt CheckEval, we identified 7 aspects and selected relevant sub-aspects for each. We formulated them as binary(Yes/No) questions. This resulted in a checklist-style evaluation sheet for each sub-aspect, enabling an intuitive and structured assessment of the generated texts. The prompts utilized for the writing evaluation are detailed in the Appendix 8.
@@ 4.2 Writing Evaluation
The evaluation criteria were selected based on prior linguistics research Wolfe (1997); Knoch (2011); van der Lee et al. (2019); Celikyilmaz et al. (2020); Chhun et al. (2022); Sai et al. (2022); van der Lee et al. (2021) and finalized through a review and refinement process involving 6 writing experts. Details of the evaluation criteria are in Figure 3, and the instructions and checklist used during the evaluation process are presented in Appendix 7.
5 Experimental Setting
5.1 Dataset
In this study, we constructed a Korean dataset based on real user-generated blog posts to assess the effects of the Writing Path framework. The dataset covers five common blog domains: travel, beauty, gardening, IT, and cooking. We selected 20 blog posts for each domain [^4] https://blog.naver.com/, resulting in a total of 100 seed data points.
For seed data construction, we generated metadata, including purpose, topic, keywords, and expected reader, based on the title and content of the selected blog posts. This metadata is the input to the Writing Path, helping the model understand the context of the post and generate relevant outlines and text. Using the seed data, we created a test dataset of 1,100 instances per model under evaluation. Each data point includes the outputs of each Writing Path step: an outline, additional information, an augmented outline, and the final text. With analysis experiments as well, we generated a total of 1,500 posts for each model, resulting in 4,500 instances in total. Each data point includes all the intermediate outputs of the Writing Path: metadata, outline, browsing output, augmented outline, and text. For human evaluation, we randomly sampled 10% of the outlines and texts and assessed their scores. The final texts were evaluated by human experts, and the dataset aligns the generated outputs from three models with the human scores.
This dataset is significant in two ways. First, it provides the materials to validate the impact of applying the Writing Path framework to LLMs. Second, by aligning model outputs with human evaluation scores, our dataset provides a valuable resource for studying writing evaluation methods. To the best of our knowledge, there is currently no Korean dataset that maps human scores to writing task outputs, making our contribution particularly valuable for advancing research in this area.
Figure 4: Overview of the main analysis steps within the Writing Path framework. Each step represents a different scenario evaluated in the study: meta-data only, initial outline, and augmented outline, respectively.
5.2 Model
In our experiments, we used three models: GPT-3.5-turbo555gpt-3.5-turbo, GPT-4666gpt-4-0125, and HyperCLOVA X [^7] HCX-003. For evaluation, we used GPT-4-turbo888gpt-4-turbo, We chose GPT-4-turbo as the evaluation model because of its state-of-the-art performance at the time of this study. model.
5.3 Human Evaluation
To evaluate the outlines and writings, we hired 12 crowd workers in total. Since the outlines are relatively simple and short, we employed 6 native Korean speakers to assess them. For the more detailed and rigorous writing evaluation, we recruited 6 professional writers and teachers as writing experts.
6 Experimental Results
6.1 Effectiveness of Writing Path
(a) (a) LLM Evaluation
(b) (b) Human Evaluation
binary binary binary binary binary binary binary binary
GPT-3.5 51.66 31.14 46.29 88.11 66.43 21.14 35.14 48.56
GPT-4 68.00 60.57 72.86 89.26 80.29 54.14 66.29 70.20
HyperCLOVA X 89.71 84.46 91.14 98.06 92.57 74.00 80.00 87.13
Figure 5 presents the results from both (a) LLM and (b) human evaluation using CheckEval. Both consistently show progressive improvement as more components of the Writing Path are incorporated, while the model rankings are in different order between the two evaluation methods [^9] In the LLM evaluation, GPT-4 scores higher than HyperCLOVA X, while the opposite is observed in the human evaluation. These differences may be attributed to the use of GPT-4-turbo as the evaluation model. They could be related to the self-enhancement bias discussed in Zheng et al. (2023). On the other hand, the superior performance of HyperCLOVA X in human evaluations can be attributed to the extensive use of Korean data during pre-training Kim et al. (2021). This can improve the model’s understanding of the Korean language and culture..
Specifically, The results show that using the augmented outline (aug) leads to better writing quality compared to using only metadata (meta), indicating that the quality of writing improves significantly when the full Writing Path pipeline is employed. Moreover, the results demonstrate that using the augmented outline (aug) yields better writing quality than using the initial outline (init), suggesting that the content enrichment process in the Writing Path further enhances writing quality.
We consider seven key aspects (Section 3) for evaluating the quality of writing. CheckEval’s binary responses for each aspect allow for identifying the specific factors contributing to the assessments. We found that logical fluency, coherence, consistency, and specificity significantly contribute to the improvement of text quality through the Writing Path (Figure 6).
During the evaluation of the writing quality, writing experts assigned binary overall quality ratings (1 for high quality, 0 for low quality) to the texts. We employed the Kendall tau correlation to examine the relationship between the overall binary ratings and the scores for each evaluation aspect. The analysis (Figure 7) revealed a significant correlation for all the aspects we designed. Interestingly, logical fluency, specificity, and coherence, which were found to be particularly important in determining the perceived quality of written content, are among the aspects that showed the most significant improvement through the Writing Path (Figure 6). The progressive improvement in these aspects can be attributed to the effectiveness of using outlines. The initial outline (init) helps organize information more logically and coherently compared to using only metadata (meta), while the augmented outline (aug) further enhances the consistency and richness of the content. These findings highlight the importance of using outlines in the writing process and demonstrate how their gradual enhancement leads to better-structured, more coherent, and content-rich texts, ultimately improving the overall quality of the written output.
@@ 6.1 Effectiveness of Writing Path
To further analyze the quality of the text generated through the complete Writing Path pipeline, we conducted a human evaluation based on the CheckEval framework. The results are presented in Table 1. The analysis by six writing experts showed that GPT-4 and HyperCLOVA X generally performed better than GPT-3.5 in terms of writing quality. HyperCLOVA X exhibited higher scores in specificity compared to other models, which is consistent with the findings reported in KMMLU Son et al. (2024) regarding the advantages of language-specific models. Detailed performance metrics across various domains and further LLM evaluations can be found in Appendix 56.

### v2

@@ Abstract
Large Language Models (LLMs) have impacted the writing process, enhancing productivity by collaborating with humans in content creation platforms. However, generating high-quality, user-aligned text to satisfy real-world content creation needs remains challenging. We propose WritingPath, a framework that uses explicit outlines to guide LLMs in generating goal-oriented, high-quality text. Our approach draws inspiration from structured writing planning and reasoning paths, focusing on reflecting user intentions throughout the writing process. To validate our approach in real-world scenarios, we construct a diverse dataset from unstructured blog posts to benchmark writing performance and introduce a comprehensive evaluation framework assessing the quality of outlines and generated texts. Our evaluations with various LLMs demonstrate that the WritingPath approach significantly enhances text quality according to evaluations by both LLMs and professional writers.
Navigating the Path of Writing:
Outline-guided Text Generation with Large Language Models
Yukyung Lee1,† Soonwon Ka2 Bokyung Son2 Pilsung Kang3 Jaewook Kang2 1Boston University 2NAVER AI Platform 3Seoul National University ylee5@bu.edupilsung_kang@snu.ac.kr {soonwon.ka, bo.son, jaewook.kang}@navercorp.com
Work done as a research intern at NAVER
@@ 1 Introduction
Writing is a fundamental means of structuring thoughts and conveying knowledge and personal opinions Collins and Gentner (1980). This process requires systematic planning and detailed review. Hayes (1980) describes writing as a complex problem-solving process and explores how planning and execution interact in writing. That is, writing involves more than merely generating text; it encompasses developing a proper understanding of the topic, gathering relevant subject matter, and implementing thorough structuring.
@@
Figure 1: Comparative overview of writing approaches: (A) direct generation, (B) iterative writing involving planning, editing, and explaining, and (C) WritingPath method, which starts with a consistency-focused plan, incorporates information-rich browsing, and results in an augmented, consistent, and rich outline.
@@
Recent advancements in Large Language Models (LLMs) have advanced the writing workflow, enhancing both its efficiency and productivity. One significant area of exploration is the collaborative use of LLMs in writing processes Lee et al. (2022); Mysore et al. (2023), as demonstrated by tools like Notion AI, Jasper, and Cohesive. The typical approach to incorporating LLMs involves the establishment of a writing plan and iterative improvement of interim outputs through revision Schick et al. (2023); Yang et al. (2022), as illustrated in Figure 1 (b) with a focus on utilizing the generative capabilities of LLMs to improve fluency, consistency, and grammatical accuracy. While these tools support users in creating content more efficiently, there remains room for improvement in maintaining consistent quality that accurately aligns with specific user intentions in production environments Wang et al. (2024).
@@
To address this, we propose WritingPath, a methodology designed to incorporate user intentions such as desired topic, textual flow, keyword inclusion, and search result integration into the writing process. WritingPath emphasizes the importance of systematic planning and a clear outline from the early stages of writing. Inspired by the structured writing plan of Hayes (1980) and the reasoning path of Wei et al. (2022), the WritingPath collects ideas and creates outlines that encapsulate the user’s intentions before generating the final text. Furthermore, the initial outlines are further augmented with additional information through information browsing. Such a structured approach offers enhanced control over the text generation process and improves the quality of the content produced by LLMs.
@@
We also utilize a multi-aspect writing evaluation framework to assess the intermediate and final productions from the WritingPath, offering a way to evaluate the quality of free-form text [^1] Free-form text generation focuses on creating diverse texts tailored to specific information and user intentions, unlike story generation, which develops narratives with plots and characters without relying on reference texts. Taking into account that conventional Likert scales (1-5 ratings) Clark ; Hinkin (1998) make it challenging to systematically compare and evaluate diverse writing outputs, particularly in creative tasks Chakrabarty et al. (2023), our evaluation framework aims to provide more precise and reliable assessments for the outlines and final texts. For evaluation purposes, we construct a free-form blog text dataset incorporating a wide range of writing styles and topics from real users, including Beauty, Travel, Gardening, Cooking, and IT. Using this dataset, we evaluate how well the LLM outputs reflect the user’s intentions. Applying the WritingPath to various LLMs shows significant performance gains across all evaluated models. These results validate that our approach enables the models to maintain a stronger focus on the given topic and purpose, ultimately generating higher-quality text that more accurately reflects user intentions. Furthermore, to validate real-world applicability, we applied WritingPath to a commercial writing platform for beta testing from October 2023 to March 2024. The deployment demonstrated its effectiveness in supporting real users with structured content creation across diverse writing needs.
@@ 2 Design of WritingPath
We propose WritingPath, a systematic writing process to produce consistent, rich, and well-organized text with LLMs. Inspired by human writing processes, it consists of five key steps: metadata preparation, initial outline generation, information browsing, augmented outline creation, and final text writing (Figure 2). Each step is guided by a specific prompt configuration that aligns LLM output with specific step requirements.
The core components of WritingPath are those that generate outlines as they establish a structured writing plan. Research suggests that a well-structured outline significantly impacts the quality of the written text Sun et al. (2022); Yang et al. (2022, 2023). The initial sketch is transformed into a detailed outline, including the flow, style, keywords, and relevant information from search results. This outline provides a clearer view of the final text to the LLMs. The specific steps are described as follows:
Figure 2: Main architecture of WritingPath, our proposed framework for guiding LLMs to generate high-quality text following a structured writing process. The WritingPath condenses text generation into five key steps. Inspired by human writing planning, it ensures alignment with specified writing goals.
Step 1: Prepare Meta Data The first step establishes the writing direction and target reader using metadata $m$, which includes i) purpose, ii) type, iii) style, and iv) keywords. To simulate this process, we converted human-written texts into metadata (see Section 4.1 for details of the dataset).
Step 2: Generate Title and Initial Outline The second step generates the title $t$ and initial outline $O_{\text{init}}$ based on the metadata $m$ from step 1, using the LLM function $f_{\text{llm}}$ with a prompt configuration function $\phi_{s}$. Here, $s$ indicates the step index, and for step 2, the prompt configuration is $\phi_{2}$:
$t,O_{\text{init}}=f_{\text{llm}}(\phi_{2}(m)),\vspace{-6pt}$ (1)
@@
Step 3: Browse for Information The third step enriches the text by collecting additional information and keywords to reinforce the initial outline. We use the search function $f_{\text{search}}$ with the generated title $t$ as the query to retrieve the top-1 blog document, $D_{\text{sim}}$:
$D_{\text{sim}}=f_{\text{search}}(t)\vspace{-6pt}$ (2)
In our implementation, we employ the NAVER search API [^2] https://developers.naver.com/docs/serviceapi to retrieve the top-1 document among similar blog posts. From the blog document, we extract keywords $K$ using the $f_{\text{llm}}$ with a prompt configuration $\phi_{3}$:
$K=f_{\text{llm}}(\phi_{3}(D_{\text{sim}})),\vspace{-6pt}$ (3)
The extracted keywords constitute the additional information from the search results, leading to a more specific writing plan, improving the quality of the generated text.
Step 4: Generate Augmented Outline The fourth step refines the initial outline by adding subheadings and specific details to each section based on incorporating the keywords collected from the previous step. The augmented outline $O_{\text{aug}}$ is generated using the LLM function $f_{\text{llm}}$ with a prompt configuration $\phi_{4}$ that takes the title $t$, keywords $K$, and initial outline $O_{\text{init}}$ as inputs:
@@
The resulting augmented outline, $O_{\text{aug}}$, comprises headers ($h_{i,0}$) and their corresponding subheaders ($h_{i,j}$), where $i$ denotes the header index, and $j$ indexes the subheaders. This detailed structure serves as a comprehensive writing plan, breaking down the text into manageable parts and providing clear direction for the content.
Step 5: Write the Text Finally, the text for each section $d^{i}$ is generated using the LLM function with a prompt configuration $\phi_{5}$ that takes the title $t$ and the corresponding section of the augmented outline $O_{\text{aug}}^{i}$ as inputs:
$d^{i}=f_{\text{llm}}(\phi_{5}(t,O_{\text{aug}}^{i}))\vspace{-6pt}$ (5)
@@
$D=\{d^{1},d^{2},\ldots,d^{n}\}\vspace{-6pt}$ (6)
WritingPath organically connects all steps of the writing process, employing an outline to aggregate and manage diverse information, and assists users in producing high-quality writing. The prompts utilized for the WritingPath are detailed in Figure 12 and 13.
@@ 3 Evaluation of WritingPath
Evaluating the effectiveness of WritingPath compared to existing writing support systems is challenging. Most previous studies do not directly utilize outlines in the writing process, resulting in a lack of systematic methods to assess outline quality. Even when outlines are used, evaluation relies only on human evaluation Yang et al. (2023); Zhou et al. (2023). Moreover, current approaches heavily rely on human evaluation, which poses challenges for assessing full texts Schick et al. (2023); Yang et al. (2022); Lee et al. (2023), as it requires evaluating multiple aspects of the written work. This challenge can arise in content creation workflows where scalable and consistent quality assessment helps maintain content standards.
To address these limitations, we propose an evaluation framework that combines human and automatic evaluation to assess the quality of generated outlines and final texts from multiple perspectives. This hybrid approach is designed to support real-world content creation workflows by combining systematic automated metrics with human assessment of nuanced writing aspects that require subjective judgment. The proposed method establishes clear evaluation criteria, enabling objective and reproducible validation of WritingPath’s effectiveness as a writing support system.
@@ 3.1.1 Automatic Evaluation
We adapt various metrics to evaluate the logical alignment, coherence, diversity, and repetition in outlines, following criteria established in linguistic literature Van Dijk (1977); Pitler and Nenkova (2008); Tang et al. (2019); Elazar et al. (2021). Logical alignment, assessed through NLI-based methods, ensures that headers and subheaders are logically connected. Coherence evaluates thematic uniformity across sections, while diversity measures the breadth of topics covered. Repetition is analyzed to minimize redundancy and improve information efficiency. Note that coherence and diversity exhibit a trade-off relationship; maintaining coherence while covering a wide range of topics is essential to ensure the effectiveness of the outline in guiding the writing process. Detailed evaluation definitions are available in Appendix C.2.
@@ 3.1.2 Human Evaluation
In addition to automatic evaluation metrics, we conduct a human evaluation to assess aspects of the generated outlines that are difficult to capture solely with automatic measures. These aspects include cohesion, natural flow, and redundancy. For augmented outlines, we also evaluate the usefulness of added information and overall improvement compared to the initial outline. Detailed evaluation definitions are available in Appendix C.3.
@@
3.2 Writing Evaluation
Traditional evaluation metrics such as Likert scales are not well-suited for assessing creative tasks like long story generation Chakrabarty et al. (2023). Acknowledging the need for more specific writing evaluation methods, we employ CheckEval Lee et al. (2024) to assess writing quality [^3] Lee et al. (2024) reports a 0.65 spearman correlation between human and LLM evaluations for dialogue, which surpasses G-Eval Liu et al. (2023). This demonstrates CheckEval’s potential as a reliable method for evaluating model-generated text quality.. CheckEval decomposes the evaluation aspects into more granular sub-questions, forming a detailed checklist. These aspect-based checklists can make performance evaluations by either humans or LLMs more fine-grained. Moreover, by explicitly capturing the evaluator’s reasoning behind each rating, this approach enhances the explainability of the evaluation process. To adapt CheckEval, we identified 7 aspects and selected relevant sub-aspects for each. We formulated them as binary (Yes/No) questions. This resulted in a checklist-style evaluation sheet for each sub-aspect, enabling an intuitive and structured assessment of the generated texts. The prompts utilized for the writing evaluation are detailed in Figure 11. The evaluation criteria were selected based on prior linguistics research Wolfe (1997); Knoch (2011); van der Lee et al. (2019); Celikyilmaz et al. (2020); Chhun et al. (2022); Sai et al. (2022); van der Lee et al. (2021) and finalized through a review and refinement process involving 6 writing experts. Details of the evaluation criteria are in Figure 3, and the instructions and checklist used during the evaluation process are presented in Table 7.
4 Experimental Setting
4.1 Dataset
In this study, we constructed a Korean dataset based on real user-written blog posts to assess the effects of the WritingPath in real content creation scenarios. The dataset covers five domains frequently handled in content creation:: travel, beauty, gardening, IT, and cooking. We created a total of 1,500 posts for each model, resulting in 4,500 instances in total. For human evaluation, we randomly sampled 10% of the outlines and texts and assessed their scores. Final texts were evaluated by human experts, aligning model outputs with professional quailty standards. Details of the dataset are in Appendix B.
Figure 4: Overview of the main analysis steps in the WritingPath framework, covering meta-data only, initial outline, and augmented outline scenarios, respectively.
4.2 Model
We conducted experiments using three models: GPT-3.5-turbo Brown et al. (2020), GPT-4 Achiam et al. (2023), and HyperCLOVA X Yoo et al. (2024)444gpt-3.5-turbo, gpt-4-0125, HCX-003. For evaluation, we used GPT-4-turbo555gpt-4-turbo; we chose GPT-4-turbo as the evaluation model because of its best performance at the time of this study.. Additionally, we attempted to adapt the WritingPath approach to open-source models, including Llama2, Orion, and KoAlpaca. However, their outputs did not meet the quality standards necessary for fair comparison, and they were excluded from our analysis.
4.3 Human Evaluation
We conducted two separate human evaluation processes, involving a total of 12 carefully selected evaluators. For outline evaluations, which are relatively simple and short, we employed 6 native Korean speakers with experience in LLM. For the more detailed and rigorous writing evaluation, we recruited 6 professional writers and teachers as writing experts, each with over 10 years of expertise in Korean writing.
5 Experimental Results
5.1 Effectiveness of WritingPath
(a) (a) LLM Evaluation
(b) (b) Human Evaluation
Figure 5 shows results from both (a) LLM and (b) human evaluation using CheckEval. Both consistently show progressive improvement as more components of the WritingPath are incorporated, while the model rankings are in different order between the two evaluation methods [^6] In the LLM evaluation, GPT-4 outperforms HyperCLOVA X, whereas the opposite trend is observed in human evaluations. These differences may be due to the use of GPT-4-turbo as the evaluation model and the self-enhancement bias discussed in Zheng et al. (2023).. Specifically, The results show that using the augmented outline (aug) leads to better writing quality compared to using only metadata (meta), indicating that the quality of writing improves significantly when the full WritingPath pipeline is employed. Furthermore, the augmented outline (aug) outperforms the initial outline (init), indicating that the content enrichment process further enhances writing quality. For a comprehensive analysis of writing quality, including human evaluation results for the final text across models, detailed improvement of text quality through the WritingPath, and Kendall tau correlations between various writing aspects and overall text quality, see Appendix D.